Consider the charges π‘ž1=2.5 πœ‡πΆ and π‘ž2=-2 πœ‡πΆ located at (π‘₯1,𝑦1,𝑧1)=(1,βˆ’2,4) π‘š and (π‘₯2,𝑦2,𝑧2)=(2,-1, 3) π‘š, respectively. Assume the charges are in the air.

a.) Determine the force exerted by charge π‘ž2Β on π‘ž1

b.) Calculate the magnitude of the force on part a.

c.) Find the electric field at (π‘₯1,𝑦1,𝑧1) induced by charge π‘ž2

d.) Determine the force exerted on charge π‘ž1Β due to the electric field of charge π‘ž2.


1
Expert's answer
2021-03-23T08:07:51-0400

a) Determine the force exerted by charge π‘ž2Β on π‘ž1:


"F=k\frac{q_1q_2}{r^2}"


b) Calculate the magnitude of the force on part a.


"F=9·10^9\frac{2.5·10^{-6}·2·10^{-6}}{(2-1)^2+(-1-(-2))^2+(3-4)^2}=0.015\text{ N}."


c) Find the electric field at (π‘₯1,𝑦1,𝑧1) induced by charge π‘ž2

"E=\frac{F}{q_1}=k\frac{q_1}{r^2}=7500\text{ V/m}."


d) Determine the force exerted on charge π‘ž1Β due to the electric field of charge π‘ž2.


"F=Eq_1=k\frac{q_1q_2}{r^2}=0.015\text{ N}."